From the given table calculate TR, MR and AR. Highlight the relationship between Total revenue and Marginal Revenue

Quantity. Price
50. 200
60. 150
70. 100
80. 50
90. 10
1
Expert's answer
2021-08-26T19:02:57-0400

TR = P×Q,

MR = (TR2 - TR1)/(Q2 - Q1),

AR = TR/Q = P.

Q P TR MR AR

50 200 10,000 -- 200

60 150 9,000 -100 150

70 100 7,000 -200 100

80 50 4,000 -300 50

90 10 900 -310 10